OpenStudy (anonymous):

Expand: (x-4)^2 A. x^2 - 16 B. x^2 - 4x + 16 C. x^2 - 8x - 16 D. x^2 + 8x - 16 E. none of these. I'm getting x^2 - 8x + 16 so E?

OpenStudy (vjt):

didn't you just ask this question?

OpenStudy (luigi0210):

Have you tried FOIL?

OpenStudy (anonymous):

@VJT you didn't explain how you got it, just answered. and @Luigi0210 yes that's how i got my answer. i did it 3 times now and i'm getting the same answer.

OpenStudy (vjt):

ok, sorry i search google and saw an explanation for it on there and just told you the answer.

OpenStudy (luigi0210):

Oh, I didn't see your answer at the bottom, sorry. But yes, the answer you are getting is correct :)

OpenStudy (vjt):

here. (x-4)^2 = (x-4)(x-4) First: x * x = x^2 Outer: x *-4 = -4x Inner: x *-4 = -4x Last: -4*-4 =16 (x-4)^2 = x^2 - 8x + 16 https://answers.yahoo.com/question/index?qid=20110920201239AAd19sy

OpenStudy (anonymous):

that's okay. so i'm getting the correct answer then. thank you both!